OpenStudy (anonymous):

If X = 5 and Y = 3, what does Z equal? 2X + 2Z = 10Y

OpenStudy (anonymous):

do you know how to substitute? 2X + 2Z = 10Y

OpenStudy (anonymous):

its actually quite easy X = 5 and Y = 3 2X + 2Z = 10Y 2 (5 ) + 2Z = 10 (3)

OpenStudy (anonymous):

try this

OpenStudy (anonymous):

then simplify

OpenStudy (anonymous):

solve for Z

OpenStudy (anonymous):

2(5)+2Z=10(3).....10+2Z=30......subract 10 to get Z alone......2Z=20.....divide by 2 to get Z!.....20/2=10! Z=10

OpenStudy (anonymous):

do this piece by piece